OpenStudy (anonymous):

What is the slope and y-intercept of x+5y=20 ?

OpenStudy (anonymous):

1/5

OpenStudy (anonymous):

for y intercept set x to zero so you have 0 + 5y=20. solve for y which would equal 4. therefore, the y intercept is (0,4) and for slope solve for y. so you would have 5y=-x+20 then divide by 5 and you would get y = -1/5 x + 4 the slope is -1/5 and the y intercept is (0,4)

OpenStudy (anonymous):

my bad, it is -1/5 :-)

OpenStudy (anonymous):

Oh ok thank you